# Training and Evaluating with Human Label Variation: An Empirical Study

### v1

@@ 5 Evaluation of HLV Training Methods
We now investigate if there is an HLV training method that consistently performs well across all datasets and evaluation metrics.
@@ 5.1 Experimental Setup
We use 5 public datasets with disaggregated annotations that have been used in previous HLV studies, four of which are in English and one in Arabic. Table 4 shows that the datasets cover both crowd (untrained) and expert (trained) annotators, as well as numbers of examples and annotations, and the language of the dataset. For HS-Brexit, MD-Agreement, and ArMIS, we use the same train–test splits as Leonardelli et al. (2023). For both ChaosNLI and MFRC, we use 10-fold cross-validation with 10% of the training portion used as a development set. For ChaosNLI, we include only the SNLI and MNLI subsets (each has 1.5K examples) because of their standard NLI setup of one premise and one hypothesis.
Text Annotation Game dataset
We also include a dataset called TAG which consists of English texts describing legal problems written by non-expert legal help seekers. We have access to this dataset because of our collaboration with Justice Connect, [^7] https://justiceconnect.org.au an Australian public benevolent institution [^8] As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile providing free legal assistance to laypeople facing legal problems. Each text in the dataset was annotated by an average of 5.5 practising lawyers, who each selected one or more out of the 33 areas of law [^9] Including a special label “Not a legal issue”. that applied to the problem. Thus, it is a multilabel classification task. Example areas of law include Neighbourhood disputes and Housing and residential tenancies. Average inter-annotator agreement ($\alpha$) over the areas of law is $0.454$, which is modest. [^10] We compute this agreement on random 10% selection of the data due to the high computational cost. This is because lawyers often have different interpretations of the same problem due to their different legal specialisations and years of experience. However, these different interpretations are all valid as human label variation because the lawyers are subject matter experts. The dataset has a total of 11K texts collected between July 2020 and early December 2023. We randomly split the dataset 8:1:1 for the training, development, and test sets respectively. While it cannot be released publicly because it is based on real-world confidential legal requests from help-seekers, we include the TAG dataset due to its importance for the meta-evaluation later in Section 6.
@@
Figure 2: Performance difference (delta) with mean MV performance on HS-Brexit, MD-Agreement, the SNLI portion of ChaosNLI (C-SNLI), MFRC, and TAG datasets. Both ArMIS and the MNLI portion of ChaosNLI are omitted because the trend is similar (their results are reported in Figure 4 in the appendix). BERT denotes either TwHIN-BERT or RoBERTa, depending on the dataset. C-SNLI doesn’t have annotator identity information, so AE, AEh, AR, ARh are inapplicable. AL is omitted from MFRC and TAG because of its non-trivial application to the multilabel case. SmF1 with TwHIN-BERT predicts zero for all test instances on HS-Brexit, while AEh with RoBERTa does so on several classes in TAG, so their entropy correlations are undefined. The methods are sorted by their mean ranking across datasets, models, and metrics.
@@ 5.2 Results
We compute the difference between the performance of a single run of an HLV training method and the mean performance of MV. We then report the mean differences across runs in Figure 2 (performance of MV is reported in Table 7 in the appendix). We make the following observations.
First, HLV training methods have different effectiveness compared to MV, with some methods performing very poorly. For example, on the SNLI portion of ChaosNLI, LA-min consistently performs worse than MV. This finding suggests the importance of choosing the right training method when embracing HLV.
@@
Third, ReL and SL are competitive with each other. However, on both HS-Brexit and the SNLI portion of ChaosNLI, ReL outperforms SL with LLaMA. This finding suggests that ReL is better suited to large language models than SL.
@@ 5.3 Discussion
That ReL and SL perform best is interesting because they are perhaps two of the most straightforward methods to incorporate HLV in model training. This finding underscores that even baseline HLV training methods have high utility. Furthermore, our finding that SL often performs best is in agreement with that of Uma et al. (2021).
That ReL performs well shows that models are capable of learning from disagreeing annotations directly, in contrast with prior work (Uma et al., 2021; Kurniawan et al., 2024). Moreover, while the results suggest that this capability isn’t limited to large language models, ReL is likely to outperform SL when such models are used. We hypothesise that this is due to the enhanced learning capacity of such models, allowing them to learn from disagreeing annotations more flexibly. We leave the investigation of this effect to future work.
Despite its strengths, ReL has substantial computational cost: because it keeps the annotations disaggregated, the size of the training data can be enormous. For example, in the ChaosNLI dataset, ReL makes the training data 100 times larger because there are exactly 100 annotations for each training instance.
@@ 6 Meta-Evaluation of HLV Evaluation Metrics
While we have determined effective HLV training methods based on performance across metrics, we haven’t yet identified which metric is the best. Intuitively, a good evaluation metric would correlate well with human preference. We now perform a meta-evaluation of the HLV evaluation metrics based on this intuition.
To obtain a reliable estimate of human preference, access to subject matter experts is essential. Therefore, when they disagree, we can be confident that the disagreement is well-founded rather than due to a misunderstanding or similar noise. In our case, we have access to qualified practising lawyers annotating the TAG dataset whose qualification is standardised and directly related to the annotation task (i.e. legal area classification). Therefore, we use the TAG dataset for this meta-evaluation despite the inability to distribute it.
@@ 6.1 Experimental Setup
For each pair of pretrained model and training method, we randomly sample an average of 28 test instances from TAG. For each instance, we (a) show the instance text and both the true and the predicted judgement distributions and (b) ask lawyers which judgement distribution [^14] Because the task is multilabel, the judgement distribution doesn’t necessarily sum to one. is more accurate. The lawyers can choose exactly one, both, or neither of them. To encourage the lawyers to form their own expectation on the judgment distribution, we first show only the instance text and ask them to annotate it with areas of law (see Figure 5 in the appendix for the annotation interface). The mean number of lawyers is 3.8, with an overall inter-annotator agreement ($\alpha$) of $0.167$. The low agreement is expected and not due to noise because the task is inherently subjective, and these annotators are practising lawyers who draw on their real-world legal experience rather than tie-breaking strategies in third-party annotation guidelines.
@@ 6.2 Results
Table 5 confirms that all metrics are positively correlated with human preference except for entropy correlation. Furthermore, it shows that only soft micro F1 and PO-JSD correlate with human judgements at a level of statistical significance. This finding suggests that both metrics are good for evaluation in the HLV context. However, none of the metrics have significant Spearman correlation coefficient.
@@
Type Metric Pearson $r$ Spearman $\rho$
Hard Micro F1 $0.442$ $0.335$
Macro F1 $0.318$ $0.271$
Soft Entropy correlation $-0.196$ $-0.338$
PO-JSD $0.692$ $0.272$
Soft micro F1 $0.611$ $0.379$
Soft macro F1 $0.291$ $0.164$
[…]
6.3 Method ranking based on human preference
Given the human annotations above, we can compute the performance of an HLV training method based on the number of times humans select the predicted judgement distribution. Table 6 shows that the rankings of the methods have some similarities to TAG results in Figure 2. For instance, both ReL and SL are ranked relatively highly, except for SL with LLaMA where it is ranked 9th. JSD is also consistently ranked highly for both RoBERTa and LLaMA, where for the former it is ranked 1st.
However, the table also shows major discrepancies. AR is consistently ranked highly, even 1st for LLaMA, but this is not at all reflected in Figure 2. LA-min is ranked 2nd and 3rd for RoBERTa and LLaMA respectively, but is often the worst method for a given dataset. That said, these discrepancies are expected because the HLV evaluation metrics only have moderate correlation with human preference (see Table 5).
6.4 Discussion
Our findings suggest that both PO-JSD and soft micro F1 are reasonably good evaluation metrics in the HLV context. Therefore, this finding provides an empirical answer to the question of the best way to perform evaluation when disagreements abound, as raised by Basile et al. (2021).
@@ 6.4 Discussion
Uma et al. (2021) pointed out that the value of PO-JSD is confined within a small range. Furthermore, PO-JSD lacks an interpretation accessible to lay people. In contrast, soft micro F1 has both an intuitive interpretation and higher Spearman correlation with human preference. Therefore, we recommend to report both metrics, but select soft micro F1 score when an accessible interpretation is important (e.g., communicating with non-technical people) or one is restricted to a single metric (e.g., in hyperparameter tuning).
@@ 7 Conclusions
We propose new training methods and evaluation metrics for human label variation (HLV). Experiments across 14 methods, 6 metrics, and 7 datasets covering both multilabel tasks and expert annotations show that training on disaggregated annotations (ReL) or soft labels (SL) performs best across evaluation metrics. Leveraging the fuzzy set theory, our proposed soft metrics have an intuitive interpretation and are reduced to standard hard metrics if there is no label variation. Our analysis shows that our proposed soft metric is weakly correlated with most existing metrics, providing unique insights into model performance. Our meta-evaluation of HLV evaluation metrics show that the proposed soft metric correlates best with human preference. However, the correlation is modest, highlighting the need for better HLV evaluation metrics.
@@
The TAG dataset is private and cannot be released publicly, which is a major limitation of our work in terms of reproducibility. This is because the data is based on real-world confidential legal requests from help-seekers. Nevertheless, we believe our work still offers valuable scientific knowledge on the investigation of HLV training methods and evaluation metrics.
Our empirical meta-evaluation is limited to a multilabel task. Thus, it is unclear if the findings extend to binary or single-label tasks. We leave this empirical meta-evaluation to future work.
Acknowledgements
This research is supported by the Australian Research Council Linkage Project LP210200917151515https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution. [^16] As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. We thank David McNamara, Yashik Anand, and Mark Pendergast also from Justice Connect for the development of the meta-evaluation annotation interface.
Appendix A Derivation of Soft F1 Scores
A.1 Fuzzy Sets
@@ A.2 Soft F1 Scores
Embracing HLV, we view judgment distributions as degrees of memberships of fuzzy sets, each corresponds to a class, over a universe of examples. Specifically, we consider $P_{ik}$ and $Q_{ik}$ as the true and the predicted degrees of membership of example $i$ in class $k$. Therefore, we can define the soft precision and recall scores using fuzzy set operations as follows (Harju and Mesaros, 2023):
@@
Appendix B Experimental Setup of Pairwise Correlation
For a given number of examples and classes, we sample $B$ matrices $\mathbf{P}$ and $\mathbf{Q}$ and compute the value of each metric on these $B$ pairs of $\mathbf{P}$ and $\mathbf{Q}$. The Pearson correlations between pairs of metrics are then computed based on these $B$ data points. We set $B=500$ for computational reasons. We report the correlation coefficients in Table 2.
Appendix C Proof of Upper-Boundedness of Soft Accuracy by PO-JSD
Theorem 1.
Proof.
$\sum_{k}\min(\pi a_{k},(1-\pi)b_{k})\\ \leq\frac{1}{2}\left(H([\pi,1-\pi])-\operatorname{JSD}_{\pi}(\mathbf{a},\mathbf{b})\right)$
for all $i$. Doubling both sides and taking the average of this inequality over $i$ give the desired result. ∎
Dataset Model Acc./Micro F1 Macro F1 Entropy corr. PO-JSD Soft acc./micro F1 Soft macro F1
@@ Appendix C Proof of Upper-Boundedness of Soft Accuracy by PO-JSD
Table 7: Mean ($\pm$ std) performance of MV measured by entropy correlation, PO-JSD, and various versions of F1 scores. C-SNLI and C-MNLI refer to the SNLI and the MNLI portions of ChaosNLI respectively. (Hard/soft) accuracy (for multiclass tasks) and (hard/soft) micro F1 score (for multilabel tasks) are reported in the same column.
(a) TwHIN-BERT on ArMIS
(b) The MNLI portion of ChaosNLI
Figure 4: Performance difference (delta) with mean MV performance on ArMIS (top) and the MNLI portion of ChaosNLI (bottom). LLaMA is omitted on ArMIS because its intended use is English while ArMIS is Arabic. C-MNLI doesn’t have annotator identity information, so AE, AEh, AR, ARh are inapplicable. The methods are sorted in the same manner as in Figure 2.
(a) Areas of law annotation
(b) Preference annotation
Figure 5: Annotation interface used to collect human preference data. Lawyers are asked to annotate areas of law for a given problem description (top) and then to select a judgment distribution is more accurate along with their certainty level (bottom).

### v2

@@ 5 Evaluation of HLV Training Methods
We now conduct an extensive study covering 6 datasets, 14 training methods, and 6 evaluation metrics across 2 language models to understand the performance landscape of HLV data.
@@ 5.1 Experimental Setup
Table 4: Public datasets used in the empirical meta-evaluation of metrics, along with the language (L), type of annotators (O) where E and C mean Experts and Crowds respectively, number of examples to the nearest thousand ($N$), average number of annotations per example ($J$), number of classes ($K$), whether annotator identity is present (A), and whether the task is multilabel classification (M).
Name Labels L O $N$ $J$ $K$ A M
HS-Brexit (Akhtar, Basile, and Patti, 2021) Hate speech EN E 1 $6.0$ 2
MD-Agreement (Leonardelli et al., 2021) Offensiveness EN C 10 $5.0$ 2
ArMIS (Almanea and Poesio, 2022) Misogyny AR E 1 $3.0$ 2
ChaosNLI (Nie, Zhou, and Bansal, 2020) NLI EN C 3 $100.0$ 3
MFRC (Trager et al., 2022) Moral sentiment EN E 18 $3.4$ 8
@@
We use 5 public datasets with disaggregated annotations that have been used in previous HLV studies, four of which are in English and one in Arabic. Table 4 shows that the datasets cover both crowd (domain non-specialist) and expert (domain specialist) annotators, as well as the number of examples, average number of annotations, and the language of the dataset. For HS-Brexit, MD-Agreement, and ArMIS, we use the same train–test splits as Leonardelli et al. (2023). For both ChaosNLI and MFRC, we use 10-fold cross-validation with 10% of the training portion used as a development set. For ChaosNLI, we include only the SNLI and MNLI subsets (each has 1.5K examples) because of their standard NLI setup of one premise and one hypothesis.
“Text Annotation Game” (TAG) dataset
We also include a dataset called TAG which consists of English texts describing legal problems written by non-expert legal help seekers. We have access to this dataset because of our collaboration with Justice Connect, [^8] https://justiceconnect.org.au an Australian public benevolent institution [^9] As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile providing free legal assistance to laypeople facing legal problems. Each text in the dataset was annotated by an average of 5.5 practising lawyers, who each selected one or more out of the 33 areas of law [^10] Including a special label “Not a legal issue”. that applied to the problem. Thus, it is a multilabel classification task. Example areas of law include Neighbourhood disputes and Housing and residential tenancies. Average inter-annotator agreement ($\alpha$) over the areas of law is $0.454$, which is modest. [^11] We compute this agreement on random 10% selection of the data due to the high computational cost. This is because lawyers often have different interpretations of the same problem due to their different legal specialisations and years of experience. However, these different interpretations are all valid as human label variation because these registered lawyers are subject matter experts who play a crucial part in interpreting the law. The dataset has a total of 11K texts collected between July 2020 and early December 2023. We randomly split the dataset 8:1:1 for the training, development, and test sets respectively. While it cannot be released publicly because it is based on real-world confidential legal requests from help-seekers, we include the TAG dataset due to its importance for the meta-evaluation later in Section 6.
@@
Figure 2: Performance difference (delta) with mean MV performance of TwHIN-BERT on ArMIS. The methods are sorted by their mean ranking across datasets, models, and metrics.
Figure 3: Performance difference (delta) with mean MV performance on HS-Brexit and MD-Agreement. SmF1 with TwHIN-BERT predicts zero for all test instances on HS-Brexit so its entropy correlation is undefined. The methods are sorted by their mean ranking across datasets, models, and metrics.
@@
Figure 4: Performance difference (delta) with mean MV performance on ChaosNLI (both the SNLI and the MNLI portions), MFRC, and TAG datasets. ChaosNLI doesn’t have annotator identity information, so AE, AEh, AR, ARh are inapplicable. AL is omitted from MFRC and TAG because it is developed for single-label rather than multilabel tasks. AEh with RoBERTa predicts zero for all test instances on several classes in TAG, so its entropy correlation is undefined. The methods are sorted by their mean ranking across datasets, models, and metrics.
@@ 5.2 Results
We compute the difference between the performance of a single run of an HLV training method and the mean performance of MV. We then report the mean differences across runs. We present the results on ArMIS in Figure 2 (the only non-English dataset), HS-Brexit and MD-Agreement in Figure 3 (Learning with Disagreements shared task datasets), and the remaining datasets in Figure 4 (multiclass and multilabel datasets). We report performance of MV is in Table 7 in the Appendix. We make the following observations.
First, HLV training methods have different effectiveness compared to MV, with some methods performing very poorly. For example, on ArMIS and both portions of ChaosNLI, LA-min consistently performs worse than MV. This finding suggests the importance of choosing the right training method when embracing HLV.
@@
Third, ReL and SL are generally competitive with each other. However, there are two exceptions to this trend. First, on ArMIS, SL outperforms ReL. Second, on both HS-Brexit and the SNLI portion of ChaosNLI, ReL outperforms SL with LLaMA. For the former finding, however, ArMIS is an outlier dataset: it is the only Arabic dataset, has the smallest number of examples and annotations per example, and all methods generally have higher variance on this dataset. The latter finding suggests that ReL is better suited to large language models than SL.
@@ 5.3 Discussion
ReL and SL appear to perform best as a training method, and that is somewhat surprising because they are the most straightforward methods to incorporate HLV in model training. That said, our finding that SL is a strong performer is in line with what Uma et al. (2021) discovered. The results of ReL, however, stand in contrast with prior work (Uma et al., 2021; Kurniawan et al., 2024). This difference may be due to the structured prediction tasks that they considered which have an exponentially large output space. We note that Uma et al. (2021) also found that ReL performs really well in simple, multiclass classification tasks, in line with our findings.
Moreover, the results suggest that ReL is likely to outperform SL when large language models are used. A possible explanation is that with SL, the model observes the complete judgement distribution of an instance as target. Thus, there is not much flexibility as the model has to predict that distribution to minimise the training loss. In contrast, with ReL, the model only observes one annotation of the instance as target. Therefore, it has more flexibility in how it distributes the judgement distribution mass across all the labels of that instance over training steps. We hypothesise that large language models have sufficient learning capacity to benefit from this flexibility while smaller models do not. This explanation is consistent with our findings on ChaosNLI that ReL has much larger gains compared to SL. ChaosNLI has 100 annotations per instance, the largest number out of all datasets. We leave further investigation of this hypothesis to future work.
Despite its strengths, ReL has substantial computational cost: because it keeps the annotations disaggregated, the size of the training data can be enormous. For example, in the ChaosNLI dataset, ReL makes the training data 100 times larger because there are exactly 100 annotations for each training instance. That ReL performs better with LLaMA than RoBERTa on ChaosNLI exacerbates the computational inefficiency of ReL: both the data and the model must be large for it to perform optimally. Mitigating this inefficiency can be an interesting avenue for future work.
Because the training data is much larger, one may think that the superiority of ReL over SL is due to its much longer training. However, this is not the case. We find that increasing the number of training iterations of SL to match that of ReL does not improve performance.
Figure 5: Areas of law annotation interface
Figure 6: Preference annotation interface
@@ 6 Meta-Evaluation of HLV Evaluation Metrics
In the previous section, we gained an understanding about the best training methods that perform consistently well across metrics. In this section, we tackle the question of what the best evaluation metric is for HLV data. We address this question by conducting a meta-evaluation of all the evaluation metrics in question. Intuitively, a good evaluation metric should correlate well with human preference. That is, if a model has a high performance according to a certain metric, we expect human judges to prefer that model’s predictions. This forms the basis of our meta-evaluation experiment.
To obtain a reliable estimate of human preference, we rely on subject matter experts. This is so that when they disagree we can be confident that the disagreement is a result of genuine variation rather than noise. In our case, we have access to practising lawyers who are more than qualified for our task (i.e. legal area classification). As such, we use the TAG dataset to conduct our meta-evaluation experiments despite the inability to distribute it.
@@ 6.1 Experimental Setup
For each pair of pretrained model and training method, we randomly sample an average of 28 test instances from TAG. For each instance, we (a) show the instance text, the ground truth label distribution and the predicted label distribution from one of our trained models, and then (b) ask lawyers which distribution [^15] Because it is multilabel task, the distribution doesn’t sum to one. is more accurate. The lawyers can choose exactly one, both, or neither of the visualisations of the distributions. On average, each instance is annotated by 3.8 lawyers. To make sure the lawyers form their own expectation on the areas of law relevant to an instance, we first show the instance text and ask them to annotate it (Figure 5). After completing this task, we then show them the label distributions from the model and ground truth (Figure 6). We conduct several pilot studies with the lawyers to refine the interface design to ensure the task is clear to our target annotators.
@@
It is worth noting that the meta-evaluation task inherently has HLV as there is not a single, objective ground truth. Every lawyer has their own preference, and our goal is to capture the collective preference of the lawyers (i.e. the proportion of lawyers who prefer the predicted distributions). To minimise the risk of lawyers figuring out which of the two judgement distribution options is the ground truth distribution, we randomise the options and the order of the instances given to the lawyers. To filter out low quality predictions, we only show areas of law whose judgement exceeds a threshold of $0.1$. [^16] We arrive at this threshold by choosing a value that results in similar numbers of labels whose judgement exceeds the value in the predicted and the ground truth distributions.
@@ 6.2 Results
Table 5 confirms that all metrics are positively correlated with human preference except for entropy correlation. Furthermore, it shows that only soft micro F1 and PO-JSD correlate with human judgements at a level of statistical significance. This finding suggests that both metrics are good for evaluation in the HLV context.
6.3 Method ranking based on human preference
Given the human annotations above, we can compute the performance of an HLV training method based on the number of times humans select the predicted judgement distribution. Table 6 shows that the rankings of the methods have some similarities to TAG results in Figures 2, 3 and 4. For instance, ReL, JSD and SL are generally strong performers (exception: SL with LLaMA). That said, we also see some discrepancies. For example, AR and LA-min come out on top here but this is not reflected in Figures 2, 3 and 4. We contend, however, that the meta-evaluation study is ultimately based on one dataset and so we should interpret these results cautiously.
@@ 6.3 Method ranking based on human preference
Type Metric Pearson $r$
Hard Micro F1 $0.442$
Macro F1 $0.318$
Soft Entropy correlation $-0.196$
PO-JSD $0.692$
Soft micro F1 $0.611$
Soft macro F1 $0.291$
@@
6.4 Discussion
Putting all the results together, PO-JSD and soft micro F1 are arguably some of the best metrics for HLV data. That said, Uma et al. (2021) argued that the value of PO-JSD is typically confined within a small range and lacks an intuitive interpretation. In contrast, soft micro F1 is more interpretable. Therefore, our general recommendation is to report both metrics, but focus on the soft micro F1 score when an accessible interpretation is important (e.g., communicating with non-technical people) or one is restricted to a single metric (e.g., in hyperparameter tuning).
@@ 7 Conclusions
We propose new evaluation metrics for evaluating model predictions with human label variation (HLV). Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. Therefore, our proposed metrics have intuitive interpretations and reduce to standard hard metrics if there is no label variation. We also generalise both our proposed metrics and some of existing ones to multilabel classification settings. We mathematically prove that our proposed soft accuracy metric is upper-bounded by an existing metric based on the Jensen-Shannon divergence.
Because the soft metrics are differentiable, we propose 3 new training methods for HLV using the metrics as the training objective. Additionally, we propose 2 new training methods that aggregate losses over annotations of the same instance. We test our proposed methods on 6 datasets spanning binary, multiclass, and multilabel classification tasks, as well as crowd and expert annotators. We evaluate across 2 language models and using a total of 6 HLV evaluation metrics including both existing and proposed ones. We compare against 9 existing HLV training methods and find that training on disaggregated annotations (ReL) or soft labels (SL) performs best across evaluation metrics, and that the former is likely to perform better than the latter when both the number of annotations per instance is large and a large language model is used. Our meta-evaluation of these metrics shows that our proposed soft metric correlates best with human preference. We recommend further research to include our proposed soft metric when reporting model performance in the HLV context.
@@
The TAG dataset is private and cannot be released publicly, which is a limitation of our work in terms of reproducibility. This is because the data is based on real-world confidential legal requests from help-seekers. More importantly, their safety and privacy is of high concern: some cases are so unique that even if you were to anonymise the cases, re-identfication may still be possible. Nevertheless, we believe our work still offers valuable scientific knowledge on the investigation of HLV training methods and evaluation metrics.
Our empirical meta-evaluation is limited to a single multilabel classification dataset. Thus, it is unclear if the findings extend to other datasets and binary or single-label tasks. Future work could expand on this limitation by experimenting with a different or diverse selection of datasets.
[…]
Derivation of Soft F1 Scores
.1 Fuzzy Sets
@@
Embracing HLV, we view judgement distributions as degrees of memberships of fuzzy sets, each corresponds to a class, over a universe of examples. Specifically, we consider $P_{ik}$ and $Q_{ik}$ as the true and the predicted degrees of membership of example $i$ in class $k$. Therefore, we can define the soft precision and recall scores using fuzzy set operations as follows (Harju and Mesaros, 2023):
@@
[…]
Performance of MV
Table 7 reports the performance of MV across datasets and evaluation metrics.
Table 7: Mean ($\pm$ std) performance of MV measured by entropy correlation, PO-JSD, and various versions of F1 scores. MD-Agr refers to MD-Agreement. C-SNLI and C-MNLI refer to the SNLI and the MNLI portions of ChaosNLI respectively. Hard (resp. soft) accuracy scores (for multiclass tasks) are reported in the hard (resp. soft) micro F1 score column.
Dataset Model Micro F1 Macro F1 Entropy corr. PO-JSD Soft micro F1 Soft macro F1
@@
Acknowledgements.
This research is supported by the Australian Research Council Linkage Project LP210200917181818https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution. [^19] As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. We thank David McNamara, Yashik Anand, and Mark Pendergast also from Justice Connect for the development of the meta-evaluation annotation interface. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.